OpenStudy (anonymous):

Y=(cubedroot(x)-1) e^x Differentiate the following

OpenStudy (anonymous):

\[f(x)=\sqrt[3]{x-1}e^x\]?

OpenStudy (anonymous):

product rule problem. \[(fg)'=f'g+g'f\] here f\ \[f(x)=\sqrt[3]{x-1},f'(x)=\frac{1}{3(x-1)^{\frac{2}{3}}}\] \[g(x)=e^x,g'(x)=e^x\] use the formula

OpenStudy (anonymous):

Yes